The measure of an angle is 28 degrees greater than it's compliment. Find the measure of each angle.
Found 2 solutions by matthew_sessoms, clhudson:
Answer by matthew_sessoms(39)

Complementary angles = 90 degrees


Lets set up the equation just like you said it.


The measure of an angle--90
is =
28 degrees greater than it's compliment (let the compliment by x)--28+x
And, of course, the compliment itself--x


Now, combine these and we get...
90 = (28 + x) + x
x = 31


The compliment angle is 31.
Now, plug 31 in for x for (28 + x), and we get 59


So the 2 angles have a measure of 31 & 59


CHECK
28 + the compliment angle + the compliment angle = 90
28 + 31 + 31 = 90
MS

Answer by clhudson(2)

let x = the angle
x-28= complement the sum of complementary angles equal 90 degrees
therefore x+(x-28)=90
2x-28=90
2x=118
x=59 angle is 59, complement is (59-28)=31